Mr and Mrs Bell have twin daughters and a son. Mr Bell is four years older than Mrs bEll. Mrs. Bell is three times as old as their twin daughters. THeir son is seven years younger than the twins. The sum of the five ages is 150. If x represents the age of the twins write an equation and solve to find the age of each family member
Answer by FrankM(1040)

Mr = Mrs+4
Mrs=3X
S=X-7
Mr+Mrs+2X+S=150
(3X+4)+3X+2X+(X-7)=150
9X-3=150
9X=153
X=17
Mrs = 51
Mr=55
S= 10